You rode your bike 24 miles and traveled at 16 miles per hour. How long did you take?
Answer by Alan3354(30993)

24/16 = 1.5 hours
-------------
You want to solve for time, "how long" is time.
24 miles has no time
16 miles/hour has time, hours in the denominator
16 miles/hr --> 1/16 hours/mile
-----------
Now you need to eliminate the miles in the denominator.
Multiply by the 24 miles, the miles units cancel.
24 miles * (1/16) hour/mile = 24/16 hours
= 1.5 hours.
-------------
Make use of the units to figure out what to do, and to check that you did the right thing.